OpenStudy (anonymous):

I'm *pretty* sure the answer here is D Which expression shows the value of cot x°? hypotenuse ÷ adjacent adjacent ÷ opposite hypotenuse ÷ opposite adjacent ÷ hypotenuse

myininaya (myininaya):

nope nope not D

myininaya (myininaya):

Unless you meant cos(x)

OpenStudy (anonymous):

How different is Cos and Cot? lol

myininaya (myininaya):

rebecca honestly there is no way for me to help you with this problem other than giving you the trig ratios which will give the answer

OpenStudy (anonymous):

No it was cot xº

OpenStudy (anonymous):

This was the problem. It was the way it was given to me. If you don't mind showing me, I'm more than willing to learn :)

OpenStudy (anonymous):

cot x = adjacent /hypotenuse option d and m sure about it

OpenStudy (anonymous):

Myin just said not D :/

myininaya (myininaya):

\[\sin(x)=\frac{opp}{hyp} , \cos(x)=\frac{adj}{hyp} , \tan(x)=\frac{opp}{adj}\] \[\csc(x)=\frac{hyp}{opp}, \sec(x)=\frac{hyp}{adj}, \cot(x)=\frac{adj}{opp}\]

OpenStudy (anonymous):

*takes notes* yes, yes.

myininaya (myininaya):